Find the value of following equation tan100 +tan125+tan 100 tan125

Grade:11

1 Answers

Aditya Gupta
4 years ago
hello tanisha.

Tan ( 100° + 125° ) = tan 225°  (but tan 225° = tan ( 180° + 45° ) = tan 45° = 1 )
 ( tan 100° + tan 125° ) / ( 1 - tan 100° tan 125° ) = 1 
 tan 100° + tan 125° = 1 - tan 100° tan 125° 
 tan 100° + tan 125° + tan 100° tan 125° = 1
kindly approve :)